Two factory plants r making tv panels...plant A produced 9000 panels...7% from plant A were defective...2% from plant B were defective...how many panels did plant B produce if the overall percentage of defective panels from both plants was 5%...this is how I set it up.....
9000 panels x 7% defective = 630 defective from plant A...X is the number of panels from plant B...2% defective .02X........so, 630+0.02X / 9000+X= 0.05.....630 +0.02X = 0.05(9000 + X)...630+0.02X= 0.05(9000) + 0.05 (X)....630 + 0.02X = 450+0.05X....630 + 0.02X - 0.05X = 450.....0.02X - 0.05X= 450 - 630.....0.03X = -180 ????

Answer by greenestamps(13200)



Fine until the last step....

0.02X - 0.05X = 450 - 630 (okay)
0.03X = -180 (not okay)

0.02X-0.05X = -0.03X, not 0.03X

Out of carelessness, or whatever else, it is easier to make mistakes working with negative numbers than with positive numbers. So I would do things a bit differently earlier in your work:

630+0.02X /(9000+X) = 0.05
630 +0.02X = 0.05(9000 + X)
630+0.02X = 0.05(9000) + 0.05 (X)
630 + 0.02X = 450+0.05X

At this point, your next step was
0.02X-0.05X = 450-630

I would move terms differently so that the "X" term is positive:

630-450 = 0.05X-0.02X
180 = 0.03X
...